Wallis Company manufactures only one product and uses a standard cost system. The company uses a predetermined plantwide overhead rate that relies on direct labor-hours as the allocation base. All of the companys manufacturing overhead costs are fixed-it does not incur any variable manufacturing overhead costs. The predetermined overhead rate is based on a cost formula that estimated $2,893,000 of fixed manufacturing overhead for an estimated allocation base of 289,300 direct labor-hours. Wallis does not maintain any beginning or ending work in process inventory The companys beginning balance sheet is as follows Wa11is Company Balance Sheet (dollars in thousands) Assets Cash Raw materials inventory Finished goods inventory Property, plant, and equipment, net Total assets Liabilities and Equity Retained earnings Total liabilities and equity $ 830 28:0 400 9.800 $11,310 $11.310 $11,310 The companys standard cost card for its only product is as follows: Standard Quantity or Bours Standard Price or Rate 2 pounds 32.60 per pound $65.20 Standard Cost Inputs Direct materials DLreet labor Pixed manufacturing overhead otal standard cost per unit 3.00 hour 14.00 per hour 3.00 hours S 10.00 per hour 42.00 30.00 $137.20 During the year Wallis completed the following transactions
media%2F98f%2F98fd9504-583e-41b0-8b16-22
Answer #1

1. Calculation of Variances :

Direct Material Variances:

Direce Material Price Variance = (SP-AP)*AQ

                                             = (30.80-29.90)*216,000 = 194,400 Fav

Direct Material Quantity variance = SP*(SQ-AQ)

                                               = 30.80*(190,800-216,000) = 776,160 Adv

*** Cal of SQ = 95,400*2 = 190,800

Total DM variance = (AP*AQ)-(SP*SQ)

                          = (29.90*216,000)-(30.8*190,800) = 581,760 Adv

Direct Labour Variances:

Price Variance = AH*(SR-AR)

                     = 245,800*(14-16) = 491,600 Adv

Quantity Variance = SR*(SH-AH)

                          = 14*(286,200-245,800) = 565,600 Fav

*** Cal of SH = 95,400*3 hr/unit = 286,200 hrs

Total DL Variance =( SR*SH)-(AR*AH)

                          (286,200*14)-(245,800*16) = 74,000Fav

Fixed Overheards variances

   Standard Rate per hour = 2,884,000/288,400 = $10/hr

   Standard hours = 95,400*3 = 286,200 hrs

   Total Standard overheads - 286,200*10=$2,862,000

   Actual overheads = 2,742,000

   Variance = 2,862,000-2,742,000 = 120,000 Adv

2. Income Statement (Actual Cost):

Sales 1,57,08,000.00
Less : Total Cost of Goods sold
Material (216,000/95,400*92,400)*29.9      62,55,305.66
Labour (245,800/95,400*92,400)*16      38,09,127.04
Fixed overheads      27,42,000.00
Selling Overheads      21,22,000.00
Total COGS 1,49,28,432.70
Net Profit                                     7,79,567.30

Please give your feedback and ask if you have any questions.